ionisation potential of hydrogen atom is 13.6 ev hydrogen atom on the ground state rarely excited by monochromatic radiation of 12.1 ev the special line emitted by a hydrogen atom according to bohr theory will be
A.1
B.2
C.3
D.4

Grade:12th pass

1 Answers

Vikas TU
3 years ago
Ionization potential
of hydrogen atom = 13.6 V
Hydrogen atoms in the ground state are excited by monochromatic radiation.
photon energy = 12.1 eV.
To find :
The spectral lines emitted by hydrogen atoms according to Bohr's =?
We know that
Ionization energy = -13.6 eV
Energy of electron in the excited state will be = ionization energy+ photon energy
= -13.6+12.1
= -1.5 eV.
Also, energy of electron in nth orbit is
E = -13.6/ n^2
n^2 = -13. 6/-1.5
n^2 = 9
n = 3
Since n = 3, the possible number of spectral lines will be transited from third to second, second to first and third to first.
Hence, the number of spectral lines are 3.
Hence,
(c) Three is the correct answer.
Note : ionization potential is the minimum required energy to remove and electron from its isolated gaseous state.